OpenStudy (anonymous):

HELP!!!!! A turntable must spin at 33.7 rpm (3.53 rad/s) to play an old-fashioned vinyl record. How much torque must the motor deliver if the turntable is to reach its final angular speed in 3.0 revolutions, starting from rest? The turntable is a uniform disk of diameter 30.5 cm and mass 0.25 kg.

OpenStudy (anonymous):

Hint: Torque = Change in Angular Momemtum ie \[L=mr^2\omega\]

OpenStudy (anonymous):

Ok, but what is L?

OpenStudy (anonymous):

starting from rest: \[\omega _{f}^2 =2 \alpha \theta \] find alpha calculate I, find torque from: \[\tau = I \alpha \]

OpenStudy (anonymous):

Thank you Algebraic!